Find the equation of the line that.contains the.point 9, -19 and has slope -2
Answer by ankor@dixie-net.com(22740)

Find the equation of the line that.contains the point 9, -19 and has slope -2
Using the point slope equation; y - y1 = m(x - x1), where
x1 = 9
y1 = -19
m = -2
y - (-19) = -2(x - 9)
y + 19 = -2x + 18
y = -2x + 18 - 19
y = -2x - 1 is the equation